OpenStudy (anonymous):

Is this equation true or false? 4(-5)=(-25) Answer: false?

OpenStudy (anonymous):

Falsely false? ;)

OpenStudy (texaschic101):

a positive times a negative is a negative...so the sign is right, but 4 x 5 = 20, not 25...it is false

OpenStudy (anonymous):

Hehe :)

OpenStudy (texaschic101):

falsely false...lol...good one d3Xter :)

OpenStudy (anonymous):

thank you

OpenStudy (texaschic101):

anytime :)

OpenStudy (anonymous):

thx texaschic101 ! :)

OpenStudy (texaschic101):

you crack me up sometimes :)

OpenStudy (anonymous):

can yous help me out with another one?

OpenStudy (texaschic101):

sure...what ya got

OpenStudy (anonymous):

solve the equation. 10(8-2)= Answer: 60? I did what was in () first then multiplied 10 and 6 and got 60

OpenStudy (texaschic101):

you are correct....you do what is in parenthesis first, then multiply.

OpenStudy (texaschic101):

or you could have distributes... 10(8 - 2) = 80 - 20 = 60 either way, you will get the same answer

OpenStudy (anonymous):

do you know this? 25/5 + 15/3 =

OpenStudy (anonymous):

im not the best with fractions

OpenStudy (texaschic101):

to add/subtract fractions, we need a common denominator. A number that 5 and 3 both go into evenly. That number would be 15. OH....wait...there is an easier way..... 25/5 = 5 and 15/3 = 5.....added together = 10

OpenStudy (anonymous):

so the answer is 10?

OpenStudy (texaschic101):

yup :)

OpenStudy (anonymous):

Thankyou do you mind helping me up with a couple more?

OpenStudy (texaschic101):

I can try :)

OpenStudy (anonymous):

2+ 8+10/2 =

OpenStudy (texaschic101):

2 + 8 + 10/2 -- (10/2 reduces to 5) 2 + 8 + 5 = 15

OpenStudy (texaschic101):

oh...wait...is the problem 2 + 8 + (10/2) or is it (2 + 8 + 10)/2 ??

OpenStudy (anonymous):

no its 2+ on the outside then next to it is the fraction 8+10/2

OpenStudy (texaschic101):

2 + (8 + 10/2) = 2 + (8 + 5) = 2 + 13 = 15 (remember, 10/2 = 5)

OpenStudy (anonymous):

Thank you i just need help with two more: 15/8-3 +6 =

OpenStudy (texaschic101):

(15/8) - 3 + 6 or 15/(8 - 3 + 6) or (15/8 - 3) + 6 all this makes a difference....

OpenStudy (texaschic101):

which one is it ?

OpenStudy (anonymous):

ik its not the second one and i got the same answer for the first and last one which was (7.125) only the first one was a negative and last one was a positive

OpenStudy (anonymous):

so i think im gonna go with the last choice

OpenStudy (texaschic101):

ok....your choice, your grade...lol

OpenStudy (anonymous):

ok just need help with one more: 6x 4/4-2 =

OpenStudy (anonymous):

and the (x) is a mutiplication sign

OpenStudy (texaschic101):

6 x (4/4) - 2 = 6 x 1 - 2 = 6 - 2 = 4 OR (6 x 4) / (4 - 2) = 24/2 = 12 OR ((6 x 4)/4) - 2 = 24/4 - 2 = 6 - 2 = 4 Take your pick...lol

OpenStudy (anonymous):

first one

OpenStudy (texaschic101):

ok....answer is 4